OpenStudy (anonymous):

5x^2+7x-66=0 I dont know if I got the right answer, can someone help me? I got x= (4.4,3)

OpenStudy (anonymous):

HINT: Use quadratic formula

OpenStudy (anonymous):

Use: x = \[x = (-b^{2} \pm \sqrt{b^{2}-4ac})/2a\] Where the eq is \[a x^{2} +bx + c\]

OpenStudy (anonymous):

\[5x ^{2}-15x+22x-66=0\] 5x(x-3)+22(x-3)=0 (5x+22)(x-3)=0